# Shift current in 2D Janus Transition-Metal Dichalcogenides: the role of excitons

### v1

@@ III.2 Linear and nonlinear response in MoSSe and WSSe
In Figs. 4 and 5 we can distinguish different excitonic peaks for both MoSSe and WSSe: in particular the A/B peak around 2 eV and the C peaks around the 3 eV. Notice that as we did not include spin-orbit coupling, the ground state exciton (A) and the spin-orbit split exciton (B) are degenerate. The strongest feature in both spectra are the C-excitons, which we found corresponds to the strongest enhancement of the SC For this reason we analyze them in detail.
We consider the two most intense excitations, C1 at 3.02 eV and C2 at 3.19 eV for WSSe, together with the A/B exciton at 2.33 eV, indicated with vertical lines in Fig. 5. We then projected their exciton weights on $k$-points in the full Brillouin zone $W^{\lambda}_{\bf k}=\sum_{cv}|A^{\lambda}_{cv{\bf k}}|^{2}$, top panel of Fig. 3, and on the valence/conduction bands along the band structure of WSSe, $W^{\lambda}_{c/v,{\bf k}}=\sum_{v/c}|A^{\lambda}_{cv{\bf k}}|^{2}$, bottom panel of Fig. 3. The A/B exciton originated from electron-hole transitions around K, the $C_{1}$ exciton originated from electron-hole transitions along K-$\Gamma$, the $C_{2}$ exciton from electron-hole transitions around $\Gamma$. For all three excitons, the holes are in the top valence band and the electrons in the bottom conduction band. By comparing the distribution of the $W^{\lambda}_{c/v,{\bf k}}$ with the atomic orbital projected band structures, (Supplemental Material), we can deduce the spatial localization of the electron-hole pairs contributing to each exciton. For the A/B exciton, both electrons and holes are mostly localized on the chalcogen atoms. Being the composition of the bottom conduction and top valence around K very similar, we deduce the electron and hole are localized around the same atom. For the $C_{1}$ exciton, while the electrons are mostly localized around the S atoms, the hole are partially localized on the Se atoms as well. For the $C_{2}$ exciton, holes are more localised on the tungsten and electrons more on the S atoms.
@@
Figure 4: Shift current along the $yyy$ direction for MoSSe (upper panel) compared with the linear absorption (lower panel), at the IP (gray dashed lines), IPGW (red dashed lines) and GW+BSE (blue solid lines) levels.
@@
Figure 5: Shift current along the $yyy$ direction in WSSe (upper panel) compared with the linear absorption (lower panel), at the IP (gray dashed lines), IPGW (red dashed lines) and GW+BSE (blue solid lines) levels.
@@
Next we consider the SC. Figures 5 and  4 show the SC of MoSSe and WSSe at the three levels of approximation.: In the nonlinear response function, the gap widening induced by GW corrections substantially modifies the spectral intensity by suppressing the SC conductivity. Including excitonic effects redshifts the full spectrum and restores its intensity, albeit with a significant redistribution with respect to the TD-IP. In both systems, the SC conductivity is strongly suppressed in correspondence of the A/B exciton, while it is strongly enhanced in correspondence of the C-excitons. A similar intensity redistribution has been observed in other 2D dichalcogenides in second-harmonic generationRuan et al. (2024) and sum-frequency generation.Pionteck et al. (2025) The relation between SHG and SC is also discussed in the work of Morinoto and Nagaosa in Ref. Morimoto and Nagaosa, 2016.
@@
To explain the stark difference of the SC at resonance with the A/B and C excitons, we refer to the analysis in terms of electron-hole pairs. While for the A/B exciton the holes and electrons are localized around the same atom, for the C excitons the electron and holes are localized on different sites. We thus argue that following a photo-excitation at resonance with the C excitons there is a shift of the center of electronic charge, which is the physical origin of the SC. We now highlight the key differences between the two systems, MoSSe and WSSe. Two important differences stand out as we move from one to the other: the first is the different band gaps, reflected in the shift of the SC spectra. The WSSe spectra starts at slightly higher energies than that of MoSSe. Second, in WSSe, C-excitons are more intense and better separated from higher excitons than in MoSSe. These differences will be important for selecting the best material having a response in the desired energy range in experimental devices. Finally, we found that the SC response of Janus dichalcogenides at resonance, renormalized with the effective thickness, has an intensity comparable to that of the strongest known 2D materials in the scientific literatureChan et al. (2021), making 2D-Janus an interesting building block for future devices.
@@ IV Conclusions
We studied the SC response of two-dimensional (2D) Janus dicalcogenides, MoSSe and WSSe, using a real-time approach that included correlation effects in the electronic band structures and response functions. We found that, in both systems, excitonic effects strongly enhance the SC conductivity at resonance with C-excitons, while suppresses it at resonance with the A/B exciton. As the SC originates from the shift of the center of electron charge following photo-excitations, we rationalize this result by noting that in the A/B exciton, the electrons and holes are localized around the same atomic site—thus no shift of the center of electron charge—while in the C excitons, the real-space distribution of electrons and holes differs significantly, so following a photoexcitation resonant with the C excitons, there is a shift of the center of electron charge and thus a photocurrent. To transform these materials into devices and take advantage of their significant response, the $\mathrm{C_{3V}}$ symmetry must be broken. This can be achieved by combining the materials into heterostructures with tensile strainHung et al. (2023), or by creating nanotubes along the non-zero response directionYang et al. (2025b). Furthermore, the possibility of inverting the SC direction by inverting the z-axis of Janus dichalcogenides and using their internal electric field to interact with closest layers endows these materials with extra tunability.
@@ Appendix A Numerical evaluation of the current operator
In Eq. (4) the current operator is written in term of a covariant $k$-derivative of the periodic part of the time-dependent valence bands. This derivative is numerically evaluated on a finite $k$-point grid as shown in Eq. (5). In this Appendix, we show how the formula for current is implemented in the Yambo code Sangalli et al. (2019). All operators and wave-functions that appear in the equations of motions are expanded in the Kohn-Sham (KS) basis set $|u_{{\bf k},n}\rangle$. In this basis, the periodic part of the time-dependent valence bands $|v_{{\bf k},n}\rangle$ and their covariant counterpart $|\widetilde{v}_{{\bf k}i\sigma,n}\rangle$ reads:
@@
where $M$ is the number of valence bands, and $\tilde{S}_{{\bf k},i\sigma}$ is the overlap matrix between real-time valence states. The indices $m,n$ run on the valence bands. $\{{\bf k}i\sigma\}={\bf k}+\sigma\Delta{\bf k}_{i}$ with $\sigma=\pm 1$ and $i$ labeling the displacement along the reciprocal lattice vectors in the ${\bf k}$-grid. The overlap matrix $S_{{\bf k},i\sigma}$ between time-dependent valence states reads:

### v2

@@ III.1 Atomic and electronic structure
For both MoSSe and WSSe we optimize atomic positions and unit cell at zero pressure by minimizing the enthalpy as implemented in the Quantum Espresso code.[pwscf] We obtained lattice parameters of 6.089 a.u. and 6.090 a.u. respectively for MoSSe and WSSe (lattice vectors and atomic coordinates are specified in the Supplemental Material). Starting from the optimized atomic structure we calculate the electronic bands and then the linear and nonlinear response functions. The electronic band structures are evaluated at the Kohn-Sham (KS) and G0W0 levels for both MoSSe and WSSe, results are shown in Fig. LABEL:fig:TMD_bands.
@@ III.2 Linear and nonlinear response in MoSSe and WSSe
In Figs. 4 and 5 we can distinguish different excitonic peaks for both MoSSe and WSSe: in particular the A/B peak around 2 eV and the C peaks around the 3 eV. Notice that as we did not include spin-orbit coupling, the ground state exciton (A) and the spin-orbit split exciton (B) are degenerate. The strongest features in both spectra are the C-excitons, which we found to correspond to the strongest enhancement of the SC. For this reason we analyze them in detail.
We consider the two most intense excitations, C1 at 3.02 eV and C2 at 3.19 eV for WSSe, together with the A/B exciton at 2.33 eV, indicated with vertical lines in Fig. 5. We then projected their exciton weights on $k$-points in the full Brillouin zone $W^{\lambda}_{\bf k}=\sum_{cv}|A^{\lambda}_{cv{\bf k}}|^{2}$, top panel of Fig. 3, and on the valence/conduction bands along the band structure of WSSe, $W^{\lambda}_{c/v,{\bf k}}=\sum_{v/c}|A^{\lambda}_{cv{\bf k}}|^{2}$, bottom panel of Fig. 3. The A/B exciton originated from electron-hole transitions around K, the $C_{1}$ exciton originated from electron-hole transitions along K-$\Gamma$, the $C_{2}$ exciton from electron-hole transitions around $\Gamma$. For all three excitons, the holes are in the top valence band and the electrons in the bottom conduction band.
By comparing the distribution of the $W^{\lambda}_{c/v,{\bf k}}$ with the atomic orbital projected band structures (see Supplemental Material), we can deduce the spatial localization of the electron-hole pairs contributing to each exciton. For the A/B exciton, both electrons and holes are mostly localized on the chalcogen atoms. Being the composition of the bottom conduction and top valence around K very similar, we deduce the electron and hole are localized around the same atom. For the $C_{1}$ exciton, while the electrons are mostly localized around the S atoms, the holes are partially localized on the Se atoms as well. For the $C_{2}$ exciton, holes are more localized on the tungsten and electrons more on the S atoms.
@@
Figure 4: Shift current along the yyyyyy direction for MoSSe (upper panel) compared with the linear absorption (lower panel), at the TD-IP (gray dashed lines), TD-IP@GW (red dashed lines) and TD-aGW(GW+BSE) (blue solid lines) levels.
@@
Figure 5: Shift current along the yyyyyy direction in WSSe (upper panel) compared with the linear absorption (lower panel), at the TD-IP (gray dashed lines), TD-IP@GW (red dashed lines) and TD-aGW (GW+BSE) (blue solid lines) levels.. The positions of A/B, C1 and C2 excitons analyzed in Fig. 3 are labeled with magenta letters in black boxes.
Next, we consider the SC. Symmetry analysis shows that SC in these Janus dichalcogenides has four independent tensor elements.[strasser2022nonlinear] Three of these tensor elements involve out-of-plane directions $\sigma^{(2)}_{z;xx}$$\sigma^{(2)}_{z;yy}$, $\sigma^{(2)}_{z;zz}$, while the fourth involves only in-plane components $\sigma^{(2)}_{y;yy}$. In this work, we focus specifically on the fourth component of the SC tensor for three reasons. First, it is the largest component [strasser2022nonlinear]. Second, the in-plane response is less sensitive to substrate interaction and therefore easier to measure. Third, our method, which is based on the dynamical Berry phase, is limited to periodic systems, and we are still working to include non-periodic directions in our real-time approach.
@@
Figures 5 and 4 show the SC of MoSSe and WSSe, respectively, at the three levels of approximation: TD-IP, TD-IP@GW, and TD-aGW. These correspond to independent-particle, independent-particle on top of $G_{0}W_{0}$ band structure, and GW+BSE, respectively.[attaccalite2011] In the nonlinear response function, the gap widening induced by $G_{0}W_{0}$ corrections substantially modifies the spectral intensity by suppressing the SC conductivity. Including excitonic effects redshifts the full spectrum and restores its intensity, albeit with a significant redistribution with respect to the TD-IP. In both systems, the SC conductivity is strongly suppressed in correspondence of the A/B exciton, while it is strongly enhanced in correspondence of the C-excitons. A similar intensity redistribution has been observed in other 2D dichalcogenides in second-harmonic generation[acs.nanolett.4c03434] and sum-frequency generation.[pionteck2025sum] The relation between second harmonic generation (SHG) and SC is also discussed in the work of Morimoto and Nagaosa in Ref. morimoto.
To explain the stark difference of the SC at resonance with the A/B and C excitons, we refer to the analysis in terms of electron-hole pairs. While for the A/B exciton the holes and electrons are localized around the same atom, for the C excitons the electron and holes are localized on different sites. We thus argue that following a photo-excitation at resonance with the C excitons there is a shift of the center of electronic charge, which is the physical origin of the SC.
@@
We now highlight the key differences between the two systems, MoSSe and WSSe. Two important differences stand out as we move from one to the other: the first is the different band gaps, reflected in the shift of the SC spectra. The WSSe spectra starts at slightly higher energies than that of MoSSe. Second, in WSSe, C-excitons are more intense and better separated from higher excitons than in MoSSe. These differences will be important for selecting the best material having a response in the desired energy range in experimental devices. Finally, we calculate the total short-circuit current for the two system studied here (see Sec. VI in Supplemental Material), and we found a maximum photo-current of 1.2 nA at the C resonance of the WSSe monolayer. This value is much larger than the maximum reported value for bulk BaTiO3[PhysRevLett.109.116601] and is 20% larger than the one reported for monolayer WS2.[PhysRevB.108.165418]
@@ IV Conclusions
We studied the SC response of two-dimensional (2D) Janus dichalcogenides, MoSSe and WSSe, using a real-time approach that includes quasi-particle corrections in the electronic band structures and takes into account local field effects and electron-hole interaction in the response functions. We found that, in both systems, excitonic effects strongly enhance the SC conductivity at resonance with C-excitons, while suppresses it at resonance with the A/B exciton. As the SC originates from the shift of the center of electron charge following photo-excitations, we rationalize this result by noting that in the A/B exciton, the electrons and holes are localized around the same atomic site—thus no shift of the center of electron charge—while in the C excitons, the real-space distribution of electrons and holes differs significantly, so following a photoexcitation resonant with the C excitons, there is a shift of the center of electron charge and thus a photocurrent. To harness this significant response in devices, the $\mathrm{C_{3V}}$ symmetry must be broken. This can be achieved by combining the materials into heterostructures with tensile strain[hung2023nonlinear], or by creating nanotubes along the non-zero response direction[mosse_tube]. Furthermore, the possibility of inverting the SC direction by inverting the z-axis of Janus dichalcogenides and using their internal electric field to interact with closest layers endows these materials with extra tunability.
@@ Appendix A Numerical evaluation of the current operator
In Eq. (4), the current operator is written in term of a covariant $k$-derivative of the periodic part of the time-dependent valence bands. This derivative is numerically evaluated on a finite $k$-point grid as shown in Eq. (5). In this Appendix, we show how the formula for current is implemented in the Yambo code [yambo]. All operators and wave-functions that appear in the equations of motions are expanded in the Kohn-Sham (KS) basis set $|u_{{\bf k},n}\rangle$. In this basis, the periodic part of the time-dependent valence bands $|v_{{\bf k},n}\rangle$ and their covariant counterpart $|\widetilde{v}_{{\bf k}i\sigma,n}\rangle$ reads:
@@
where $M$ is the number of valence bands, and $\tilde{S}_{{\bf k},i\sigma}$ is the overlap matrix between real-time valence states. The indices $m,n$ run on the valence bands. $\{{\bf k}i\sigma\}={\bf k}+\sigma\Delta{\bf k}_{i}$ with $\sigma=\pm 1$ and $i$ labeling the displacement along the reciprocal lattice vectors in the $\bf k$}-grid.
The overlap matrix $S_{{\bf k},i\sigma}$ between time-dependent valence states reads: